# Nonlocal pseudo-symmetries and Bäcklund transformations as $\mathcal{C}$-morphisms

### v2

@@ 2.3.2 Pseudosymmetries of differential equations
The heat equation =$\mathcal{E}=\{u_{t}-u_{xx}=0\}$, where $u=u(x,t)$, admits the classical infinitesimal symmetry (which is a particular pseudosymmetry of $\mathcal{E}$ with $\mu=0$) $Y=u\,\partial_{u}+u_{x}\partial_{u_{x}}+u_{t}\partial_{u_{t}}+\,...$ provided by the prolongation of $u\,\partial_{u}$; indeed it is readily seen that $Y(u_{t}-u_{xx})=u_{t}-u_{xx}=0\mod{\mathcal{E}}$. In this case, since $Y$ is a classical symmetry of the ambient Cartan distribution, it projects to a vector field on any finite order jet space $J^{k}(\pi)$ (where $\pi:\mathbb{R}^{2}\times\mathbb{R}\rightarrow\mathbb{R}^{2},\;(x,t,u)\mapsto(x,t)$). On the other hand, being also tangent to $\mathcal{E}$ (hence a symmetry of $\mathcal{E}$), its restriction $\bar{Y}:=\left.Y\right|_{\mathcal{E}^{(\infty)}}$ projects to a vector field on $\mathcal{E}$ as well as any of its prolongations $\mathcal{E}^{(\ell)}$. From now one we will limit ourself to domains where $u$ and its derivatives do not vanish, hence above projections of $Y$ are always nontrivial. Thus, being $\mathcal{E}$ a $7$-dimensional submanifold of $J^{2}(\pi)$, $\bar{Y}$ admits $16=\dim\mathcal{E}-$ 1 functionally independent invariants of order $2$, 8=dim($18=\dim\mathcal{E}^{(1)}-$ 1 functionally independent invariants of order $3$, and so on. These invariants can be obtained by means of prolongations from the following basic system of invariants $\{\xi^{1}=x,\,\xi^{2}=t,\,\nu^{1}=u_{x}/u,\,\nu^{2}=u_{t}/u\}$ of $\bar{Y}$, that define a $\mathcal{C}$-morphism $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$, where $\pi^{\prime}:\mathbb{R}^{2}\times\mathbb{R}^{2}\rightarrow\mathbb{R}^{2},\;(x,t,u^{1}{}^{\prime},u^{2}{}^{\prime})\mapsto(x,t)$. Indeed, non-degeneracy condition (4) is satisfied and a first prolongation provides
@@
Hence, a complete system of second order functionally independent invariants of $\bar{Y}$ is given by $\{x,\,t,\,u^{1}{}^{\prime},\,u^{2}{}^{\prime},\,u^{2}{}^{\prime}_{x},\,u^{2}{}^{\prime}_{t}\}$, since the other two invariants $u^{1}{}^{\prime}_{x},u^{1}{}^{\prime}_{t}$ satisfy the relations
$u^{1}{}^{\prime}_{x}=u^{2}{}^{\prime}-\left(u^{1}{}^{\prime}\right)^{2},\qquad u^{1}{}^{\prime}_{t}=u^{2}{}^{\prime}_{x}\qquad\mod{\{u_{xx}=u_{t}\}}.$
This means that $\mathcal{B}$ is a $\mathcal{C}$-morphism from the heat equation =$\mathcal{E}=\{u_{t}-u_{xx}=0\}$ to the system $\mathcal{Q}=\{u^{1}{}^{\prime}_{x}-u^{2}{}^{\prime}+\left(u^{1}{}^{\prime}\right)^{2}=0,\,u^{1}{}^{\prime}_{t}-u^{2}{}^{\prime}_{x}=0\}$, that one can interpret as the factorization of $\mathcal{E}$ by $Y$. Hence, $\left\{u^{1}{}^{\prime}=u_{x}/u,\,u^{2}{}^{\prime}=u_{t}/u\right\}$ transforms solutions of $\mathcal{E}$ to solutions of $\mathcal{Q}$, as well as of any of its differential consequences. In particular, since the Burgers equation =$\mathcal{E}^{\prime}=\{u^{\prime}_{t}-u^{\prime}_{xx}-2u^{\prime}u^{\prime}_{x}=0\}$ with $u^{\prime}:=u^{1}{}^{\prime}$ is a differential consequence of $\mathcal{Q}$, one obtains the Cole-Hopf transformation $u^{\prime}=u_{x}/u$ as a by-product of above transformation.
@@
The mKdV equation =$\mathcal{E}=\{u_{t}-u_{xxx}+6u^{2}u_{x}=0\}$, can be written in conservation law form as $D_{t}\left(u\right)+D_{x}\left(-u_{xx}+2u^{3}\right)=0$. Thus any $\mu=c\left(udx+(u_{xx}-2u^{3})dt\right)$, with $c\in\mathbb{R}$, is a conservation law of $\mathcal{E}$ and one can search for pseudosymmetries of the form (17), where $x_{1}=x$ and $x_{2}=t$. Here we consider the pseudosymmetry found by Sokolov in [38], obtained from the vector field $\partial_{u}$ by a pseudo-prolongation relative to $2\mu$:
@@
[…]
@@ 2.4 Differentiable coverings and Bäcklund transformations as nonlocal $\mathcal{C}$-morphisms
According to above definition (see [8]), when $\tau$ has finite fiber dimension $q$, $\tilde{\mathcal{E}}$ is locally diffeomorphic to the product ()×$\mathcal{E}^{(\infty)}\times W$, where $W\subseteq\mathbb{R}^{q}$ is an open set. Thus, locally identifying $\tilde{\mathcal{E}}$ with ()×$\mathcal{E}^{(\infty)}\times W$, $\tau$ can be locally realized as the natural projection ()×$\mathcal{E}^{(\infty)}\times W\rightarrow\mathcal{E}^{(\infty)}$. Then, using standard coordinates $v^{1},...,v^{q}$ in $\mathbb{R}^{q}$, the generators $\tilde{D}_{i}$ of the involutive distribution $\tilde{\mathcal{C}}$ on ()×$\mathcal{E}^{(\infty)}\times W$ can be locally written as $\tilde{D}_{i}=\bar{D}_{i}+\sum_{s}X_{i}^{s}\partial_{v^{s}}$ where $\bar{D}_{i}$ are the total derivatives on $\mathcal{E}^{(\infty)}$ and $X_{i}^{s}$ are smooth functions such that $\left[\tilde{D}_{i},\tilde{D}_{j}\right]=0$, i.e., $\tilde{D}_{i}\left(X_{j}^{s}\right)=\tilde{D}_{j}\left(X_{i}^{s}\right)$, in view of $[\bar{D}_{i},\bar{D}_{j}]=0$.
In this paper we will consider only the case when $\tilde{\mathcal{E}}$ is the infinite prolongation of a formally integrable system $\mathcal{Y}\subset J^{l}(\pi)$. Thus, in view of above considerations, the fiber coordinates $\{u^{1},...,u^{m}\}$ of $\pi$ will be separated in two subsets $\{z^{1},...,z^{m-q}\}$ and $\{v^{1},...,v^{q}\}$, such that =$\mathcal{E}=\left\{\mathbf{F}(x,z_{\sigma})=0\right\}$ and $\mathcal{Y}^{(\infty)}$ is the infinite prolongation of a system
@@
It is noteworthy to observe that, in view of the definition of a differentiable covering $\tau:\mathcal{Y}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$, any solution of $\mathcal{Y}$ can be projected by $\tau$ to a solution of $\mathcal{E}$.
A system $\mathcal{Y}$ like (23) will be called a differentiable extension of =$\mathcal{E}=\left\{\mathbf{F}=0\right\}$. Moreover, given a differentiable extension $\mathcal{Y}$ of $\mathcal{E}$, it is customary to call $v^{1},...,v^{q}$ the nonlocal variables of the given extension (or the corresponding differentiable covering). Also, non-local symmetries of $\mathcal{E}$ are the symmetries of a differentiable extension $\mathcal{Y}$ of $\mathcal{E}$.
@@
It is well known [31] (see also [26, 13, 30, 41] and references therein) that KdV equation =$\mathcal{E}=\{z_{t}=-6zz_{x}-z_{xxx}\}$, where $z=z(x,t)$, is the integrability condition of the system (Lax pair)
$\left\{\begin{array}[]{l}\phi_{xx}=-z\phi-\lambda\phi,\vspace{5pt}\\ \phi_{t}=2(2\lambda-z)\phi_{x}+z_{x}\phi,\vspace{5pt}\end{array}\right.$ (24)
where $\phi=\phi(x,t)$ and $\lambda$ is a real parameter usually referred to as the spectral parameter. Geometrically this means that (24) defines a differentiable covering of KdV. Indeed, by setting $v^{1}:=\phi$ and $v^{2}:=\phi_{x}$, one can rewrite (24) as a first order system
$\left(\begin{array}[]{l}v_{x}^{1}\vspace{5pt}\\ v_{x}^{2}\end{array}\right)=X\left(\begin{array}[]{l}v^{1}\vspace{5pt}\\ v^{2}\end{array}\right),\qquad\left(\begin{array}[]{l}v_{t}^{1}\vspace{5pt}\\ v_{t}^{2}\end{array}\right)=T\left(\begin{array}[]{l}v^{1}\vspace{5pt}\\ v^{2}\end{array}\right),$
with
$X=\left(\begin{array}[]{ll}0&1\vspace{5pt}\\ -z-\lambda&0\end{array}\right),\qquad T=\left(\begin{array}[]{ll}z_{x}&4\lambda-2z\vspace{5pt}\\ z_{xx}+2z^{2}-2\lambda z-4\lambda^{2}&-z_{x}\end{array}\right)$
usually referred to as a matrix Lax pair for the KdV.
@@
Thus, by introducing the fiber bundle $\mathbb{R}^{4}\rightarrow\mathbb{R}^{2}$, $(x,t,z,v^{1},v^{2})\mapsto(x,t)$, it is easy to check that the infinite prolongation of the system $\mathcal{V}$
@@
$\left\{\begin{array}[]{l}v_{x}^{1}=v^{2},\vspace{5pt}\\ v_{x}^{2}=-(z+\lambda)v^{1},\vspace{5pt}\\ v_{t}^{1}=z_{x}v^{1}+\left(4\lambda-2z\right)v^{2},\vspace{5pt}\\ v_{t}^{2}=\left(z_{xx}+2z^{2}-2\lambda z-4\lambda^{2}\right)v^{1}-z_{x}v^{2},\vspace{5pt}\\ z_{t}=-6zz_{x}-z_{xxx}\end{array}\right.$ (25)
defines a $2$-dimensional differentiable covering $\mathcal{V}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$ of KdV.
@@
It is interesting to notice here that this covering is an extension of a $1$-dimensional covering, by means of the potential of a conservation law. Indeed, by introducing $\rho:=v^{2}/v^{1}$ one can readily see that (25) reduces to
$\left\{\begin{array}[]{l}\left(\ln v^{1}\right)_{x}=\rho,\vspace{5pt}\\ \left(\ln v^{1}\right)_{t}=z_{x}+\left(4\lambda-2z\right)\rho,\vspace{5pt}\\ \rho_{x}=-\rho^{2}-z-\lambda,\vspace{5pt}\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vspace{5pt}\\ z_{t}=-6zz_{x}-z_{xxx}.\end{array}\right.$
@@
It follows that above covering extends the $1$-dimensional differentiable covering $\tau:\mathcal{Y}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$, defined by the system $\mathcal{Y}$

### v5

@@ 2.3.2 Pseudosymmetries of differential equations and factorisation
The heat equation $\mathcal{E}=\{u_{t}-u_{xx}=0\}$, where $u=u(x,t)$, admits the classical infinitesimal symmetry $Y=u\,\partial_{u}+u_{x}\partial_{u_{x}}+u_{t}\partial_{u_{t}}+\,$... (which defines a pseudosymmetry $\bar{Y}$ of $\mathcal{E}$ with $\mu=0$) provided by the prolongation of $u\,\partial_{u}$; indeed it is readily seen that $Y(u_{t}-u_{xx})=u_{t}-u_{xx}=0\mod{\mathcal{E}}$. In this case, since $Y$ is a classical symmetry of the ambient Cartan distribution, it projects to a vector field on any finite order jet space $J^{k}(\pi)$ (where $\pi:\mathbb{R}^{2}\times\mathbb{R}\rightarrow\mathbb{R}^{2},\;(x,t,u)\mapsto(x,t)$). On the other hand, being also tangent to $\mathcal{E}$ (hence a symmetry of $\mathcal{E}$), its restriction $\bar{Y}:=\left.Y\right|_{\mathcal{E}^{(\infty)}}$ projects to a vector field on $\mathcal{E}$ as well as any of its prolongations $\mathcal{E}^{(\ell)}$. From now one we will limit ourself to domains where $u$ and its derivatives do not vanish, hence above projections of $Y$ are always nontrivial. Thus, being $\mathcal{E}$ a $7$-dimensional submanifold of $J^{2}(\pi)$, $\bar{Y}$ admits $6=\dim\mathcal{E}-1$ functionally independent invariants of order $2$, $8=\dim\mathcal{E}^{(1)}-1$ functionally independent invariants of order $3$, and so on. These invariants can be obtained by means of prolongations from the following basic system of invariants $\{\xi^{1}=x,\,\xi^{2}=t,\,\nu^{1}=u_{x}/u,\,\nu^{2}=u_{t}/u\}$ of $\bar{Y}$, that define a $\mathcal{C}$-morphism $\mathcal{B}:J^{\infty}(\pi)\rightarrow J^{\infty}(\pi^{\prime})$, where $\pi^{\prime}:\mathbb{R}^{2}\times\mathbb{R}^{2}\rightarrow\mathbb{R}^{2},\;(x,t,u^{1}{}^{\prime},u^{2}{}^{\prime})\mapsto(x,t)$. Indeed, non-degeneracy condition (4) is satisfied and a first prolongation provides
@@
Thus, since the other two invariants $u^{1}{}^{\prime}_{x},u^{1}{}^{\prime}_{t}$ satisfy the relations
$u^{1}{}^{\prime}_{x}=u^{2}{}^{\prime}-\left(u^{1}{}^{\prime}\right)^{2},\qquad u^{1}{}^{\prime}_{t}=u^{2}{}^{\prime}_{x}\qquad\mod{\{u_{xx}=u_{t}\}},$
a basic system of second order functionally independent invariants of $\bar{Y}$ is given by $\{x,\,t,\,u^{1}{}^{\prime},\,u^{2}{}^{\prime},\,u^{2}{}^{\prime}_{x},\,u^{2}{}^{\prime}_{t}\}$. This means that $\mathcal{B}$ is a $\mathcal{C}$-morphism from the heat equation $\mathcal{E}=\{u_{t}-u_{xx}=0\}$ to the system $\mathcal{Q}=\{u^{1}{}^{\prime}_{x}-u^{2}{}^{\prime}+\left(u^{1}{}^{\prime}\right)^{2}=0,\,u^{1}{}^{\prime}_{t}-u^{2}{}^{\prime}_{x}=0\}$, that one can interpret as the factorisation of $\mathcal{E}$ by $\bar{Y}$. Hence, $\left\{u^{1}{}^{\prime}=u_{x}/u,\,u^{2}{}^{\prime}=u_{t}/u\right\}$ transforms solutions of $\mathcal{E}$ to solutions of $\mathcal{Q}$, as well as of any of its differential consequences. In particular, since the Burgers equation $\mathcal{E}^{\prime}=\{u^{\prime}_{t}-u^{\prime}_{xx}-2u^{\prime}u^{\prime}_{x}=0\}$ with $u^{\prime}:=u^{1}{}^{\prime}$ is a differential consequence of $\mathcal{Q}$, one obtains the Cole-Hopf transformation $u^{\prime}=u_{x}/u$ as a by-product of above transformation.
@@
The mKdV equation $\mathcal{E}=\{u_{t}-u_{xxx}+6u^{2}u_{x}=0\}$, can be written in conservation law form as $D_{t}u+D_{x}\left(-u_{xx}+2u^{3}\right)=0$. Thus any $\mu=c\left(udx+(u_{xx}-2u^{3})dt\right)$, with $c\in\mathbb{R}$, is a conservation law of $\mathcal{E}$ and one can search for pseudosymmetries of the form (17), where $x_{1}=x$ and $x_{2}=t$. Here we consider the pseudosymmetry $\bar{Y}$ found by Sokolov in [41], obtained from the vector field $\partial_{u}$ by a pseudoprolongation relative to $2\mu$:
@@
In this case, the projections of $Y$ to $J^{k}(\pi)$ (where $\pi:\mathbb{R}^{2}\times\mathbb{R}\rightarrow\mathbb{R}^{2},\;(x,t,u)\mapsto(x,t)$) are relative vector fields. The same occurs for projections of $Y$ to $\mathcal{E}$ and prolongations $\mathcal{E}^{(\ell)}$. For instance, the projection of $Y$ on $J^{1}(\pi)$ is the relative vector field $\partial_{u}+2u\partial_{u_{x}}+2(u_{xx}-2u^{3})\partial_{u_{t}}$. As a consequence, $\bar{Y}$ has only the $3$ functionally independent invariants of first order $\{\xi^{1}=x,\,\xi^{2}=t,\,\nu^{1}=u_{x}-u^{2}\}$, $3$ further functionally independent invariants of second order $\{\nu_{x}^{1}=u_{xx}-2uu_{x},\,\nu_{t}^{1}=u_{xt}-2uu_{t},\,\nu^{2}=u_{t}-2u\,\nu_{x}^{1}-2u^{2}\nu^{1}\}$ and only $3$ of the derived invariants of third order $\{\nu_{x}^{2},\nu_{y}^{2},\nu_{xx}^{1},\nu_{xy}^{1},\nu_{yy}^{1}\}$ are functionally independent since $2\nu^{1}\nu_{x}^{1}+\nu_{x}^{2}-\nu_{t}^{1}=0$, $2\nu^{1}+\nu_{xx}^{1}-\nu^{2}=0$. Then all higher order invariants are generated by $\{\xi^{i},\,\nu_{\mu}^{j}\}$ and the factorisation of $\mathcal{E}$ by $\bar{Y}$ is described by the system of differential equations $\mathcal{Q}=\{2\nu^{1}\nu_{x}^{1}+\nu_{x}^{2}-\nu_{t}^{1}=0,\,2\nu^{1}+\nu_{xx}^{1}-\nu^{2}=0\}$, that describes the lower order functional dependencies between the invariants $\{\xi^{i},\,\nu_{\mu}^{j}\}$. Hence, the map $\left\{\nu^{1}=u_{x}-u^{2},\,\nu^{2}=u_{t}-2u\,\left(u_{xx}-2uu_{x}\right)-2u^{2}\left(u_{x}-u^{2}\right)\right\}$ transforms solutions of $\mathcal{E}$ to solutions of $\mathcal{Q}$, as well as of any of its differential consequences. In particular, since the KdV equation $\mathcal{E}^{\prime}=\{u^{\prime}_{t}-u^{\prime}_{xxx}-6u^{\prime}u^{\prime}_{x}=0\}$ with $u^{\prime}:=\nu^{1}$ is a differential consequence of $\mathcal{Q}$, one obtains the Miura transformation $u^{\prime}=u_{x}-u^{2}$ as a by-product of above transformation.
@@ 2.4 Differentiable coverings and Bäcklund transformations as nonlocal $\mathcal{C}$-morphisms
According to above definition (see [8]), when $\tau$ has finite fiber dimension $q$, $\tilde{\mathcal{E}}$ is locally diffeomorphic to the product $\mathcal{E}^{(\infty)}\times W$, where $W\subseteq\mathbb{R}^{q}$ is an open set. Thus, locally identifying $\tilde{\mathcal{E}}$ with $\mathcal{E}^{(\infty)}\times W$, $\tau$ can be locally realized as the natural projection $\mathcal{E}^{(\infty)}\times W\rightarrow\mathcal{E}^{(\infty)}$. Then, using standard coordinates $v^{1},...,v^{q}$ in $\mathbb{R}^{q}$, the generators $\tilde{D}_{i}$ of the involutive distribution $\tilde{\mathcal{C}}$ on $\mathcal{E}^{(\infty)}\times W$ can be locally written as $\tilde{D}_{i}=\bar{D}_{i}+\sum_{s}X_{i}^{s}\partial_{v^{s}}$ where $\bar{D}_{i}$ are the total derivatives on $\mathcal{E}^{(\infty)}$ and $X_{i}^{s}$ are smooth functions such that $\left[\tilde{D}_{i},\tilde{D}_{j}\right]=0$, i.e., $\tilde{D}_{i}X_{j}^{s}=\tilde{D}_{j}X_{i}^{s}$, in view of $\bar{D}_{i},\bar{D}_{j}]=0$.
In this paper we will consider only the case when $\tilde{\mathcal{E}}$ is the infinite prolongation of a formally integrable system $\mathcal{V}\subset J^{l}(\pi)$. Thus, in view of above considerations, the fiber coordinates $\{u^{1},...,u^{m}\}$ of $\pi$ will be separated in two subsets $\{z^{1},...,z^{m-q}\}$ and $\{v^{1},...,v^{q}\}$, such that $\mathcal{E}=\left\{\mathbf{F}(x,z_{\sigma})=0\right\}$ and $\mathcal{V}^{(\infty)}$ is the infinite prolongation of a system
@@
It is noteworthy to observe that, in view of the definition of a differentiable covering $\tau:\mathcal{V}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$, any solution of $\mathcal{V}$ can be projected by $\tau$ to a solution of $\mathcal{E}$.
A system $\mathcal{V}$ like (24) will be called a $q$-dimensional differentiable extension of $\mathcal{E}=\left\{\mathbf{F}=0\right\}$. Moreover, given a differentiable extension $\mathcal{V}$ of $\mathcal{E}$, it is customary to call $v^{1},...,v^{q}$ the nonlocal variables of the given extension (or the corresponding differentiable covering). Also, non-local symmetries and nonlocal conservation laws of $\mathcal{E}$ are the symmetries and conservation laws of a differentiable extension $\mathcal{Y}$ of $\mathcal{E}$.
Examples of differential extensions, hence of coverings, can be easily obtained by using in (24) the first order equations for the potentials of conservation laws, when $n=2$. For instance, the equations $v_{x}=u$, $v_{t}=u_{xx}-2u^{3}$ for the potential $v$ of the conservation law $u\,dx+(u_{xx}-2u^{3})\,dt$ of mKdV (see Example 14), provide a differentiable extension of mKdV.
Other examples, that are particularly relevant in the study of Bäcklund transformations, are related to ZCRs. Indeed condition (1), for an $\ell\times\ell$ matrix-valued ZCR $\alpha=X\,dx+T\,dt$, is equivalent to the integrability condition of a linear system
$\left(\begin{array}[]{l}v_{x}^{1}\vskip 5.0pt\\ \vdots\\ v_{x}^{\ell}\end{array}\right)=X\left(\begin{array}[]{l}v^{1}\vskip 5.0pt\\ \vdots\\ v^{\ell}\end{array}\right),\qquad\left(\begin{array}[]{l}v_{t}^{1}\vskip 5.0pt\\ \vdots\\ v_{t}^{\ell}\end{array}\right)=T\left(\begin{array}[]{l}v^{1}\vskip 5.0pt\\ \vdots\\ v^{\ell}\end{array}\right).$
@@
The KdV equation $\mathcal{E}=\{z_{t}=-6zz_{x}-z_{xxx}\}$, with $z=z(x,t)$, admits the ZCR $\alpha=X\,dx+T\,dt$ given by
$X=\left(\begin{array}[]{ll}0&1\vskip 5.0pt\\ -z-\lambda&0\end{array}\right),\qquad T=\left(\begin{array}[]{ll}z_{x}&4\lambda-2z\vskip 5.0pt\\ z_{xx}+2z^{2}-2\lambda z-4\lambda^{2}&-z_{x}\end{array}\right),$
@@
with $\lambda$ a real parameter usually referred to as the spectral parameter. Thus, by introducing the fiber bundle $\mathbb{R}^{2}\times\mathbb{R}^{3}\rightarrow\mathbb{R}^{2}$, $(x,t,z,v^{1},v^{2})\mapsto(x,t)$, the infinite prolongation of the linear system
@@
$\mathcal{V}:\quad\left\{\begin{array}[]{l}v_{x}^{1}=v^{2},\vskip 5.0pt\\ v_{x}^{2}=-(z+\lambda)v^{1},\vskip 5.0pt\\ v_{t}^{1}=z_{x}v^{1}+\left(4\lambda-2z\right)v^{2},\vskip 5.0pt\\ v_{t}^{2}=\left(z_{xx}+2z^{2}-2\lambda z-4\lambda^{2}\right)v^{1}-z_{x}v^{2},\vskip 5.0pt\\ z_{t}=-6zz_{x}-z_{xxx},\end{array}\right$. (25)
defines a $2$-dimensional differentiable covering $\mathcal{V}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$ of $\mathcal{E}$. It is noteworthy to remark that, by setting $v^{1}:=\phi$ and $v^{2}:=\phi_{x}$, the above first order linear system can also be rewritten in the following equivalent form (Lax pair)
$\left\{\begin{array}[]{l}\phi_{xx}=-z\phi-\lambda\phi,\vskip 5.0pt\\ \phi_{t}=2(2\lambda-z)\phi_{x}+z_{x}\phi.\vskip 5.0pt\end{array}\right$. (26)
Indeed, it is well known [34] (see also [29, 14, 33, 44] and references therein) that $\mathcal{E}$ is the integrability condition of (26).
@@
Also, it is interesting to notice here that the considered $2$-dimensional differentiable covering, defined by (25), is an extension of a $1$-dimensional covering, by means of the equations for the potential of a conservation law. Indeed, by introducing $\rho:=v^{2}/v^{1}$ one can readily see that (25) reduces to
$\left\{\begin{array}[]{l}\left(\ln v^{1}\right)_{x}=\rho,\vskip 5.0pt\\ \left(\ln v^{1}\right)_{t}=z_{x}+\left(4\lambda-2z\right)\rho,\vskip 5.0pt\\ \rho_{x}=-\rho^{2}-z-\lambda,\vskip 5.0pt\\ \rho_{t}=-2(2\lambda-z)\rho^{2}-2z_{x}\rho+z_{xx}+2z^{2}-2\lambda z-4\lambda^{2},\vskip 5.0pt\\ z_{t}=-6zz_{x}-z_{xxx}.\end{array}\right$.
@@
Thus (25) describes an extension of the $1$-dimensional differentiable covering $\tau:\mathcal{Y}^{(\infty)}\longrightarrow\mathcal{E}^{(\infty)}$, defined by the system